# Adaptive Behavior with Stable Synapses

### v4

@@ Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated learning theories to synaptic plasticity. However, such rapid changes are not coherent with the timescales of biological synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical reconfiguration of the network involved, without changing its weights. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks, such as in pyramidal neurons thanks to input segregation and dendritic amplification. We propose a constructive approach to induce in-context learning in an architecture composed of recurrent networks with gain modulation, demonstrating abilities inaccessible to standard networks. In particular, we show that, such architecture can dynamically implement standard gradient-based by encoding weight changes in the activity of another network. We argue that, while these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the level of neural circuits. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
@@ Introduction
Main / Framework description
Dynamical adaptation replaces synaptic plasticity: A popular example
Consider a task inspired by the Wisconsin card-sorting test. Each card in a deck is defined by two features: color and shape. In each iteration, the agent receives a cue card and two guide cards—one matching the cue by color and the other by shape. The agent must select the correct guide card to receive a positive reward $R(t)=1$. Initially, the correct choice is based on color matching. However, after 10 trials, the classification rule changes, rewarding shape matching instead. A successful agent must quickly adapt its decision-making to these contextual shifts.
We consider a simple neural network that predicts reward probability for each (cue card, guide card) pair. Two populations, $z_{1}$ and $z_{2}$, encode similarity based on color and shape, respectively. A third population, $y=\phi(w_{1}z_{1}+w_{2}z_{2})$, integrates this information to predict the reward probability. Here $\phi(x)=\text{sigmoid}(x-\text{bias})$ represents the neural transfer function. The weights $w_{1}$ and $w_{2}$ combine the features to determine the prediction, and the guide card with the highest predicted reward probability is selected. The agent dynamically updates the model based on received rewards to adapt to changing contexts.
Traditionally, weights $w_{1}$ and $w_{2}$ are considered synaptic weights, and learning is attributed to synaptic plasticity. Alternatively, $w_{1}$ and $w_{2}$ can be reinterpreted as the activity of two additional populations acting as gain modulators In our model, we refer to them as "virtual weights," as they replace traditional physical synaptic weights, effectively virtualizing the learning process. These modulatory populations dynamically reconfigure the responses of the network, enabling fast, flexible adaptations without the need for synaptic changes. One of the key advantages of this virtualization is that it allows for learning rules that are not constrained by the strict locality requirements inherent to synaptic plasticity.
To support this point, in Fig. 1 we compare synaptic learning using the gradient-based, delta rule $\Delta\mathbf{w}=\eta e\cdot\mathbf{z}$ with a non-local learning rule that force constraints $w_{1}(t)+w_{2}(t)=const$. This results in the update rule $\Delta w_{j}=\eta(-1)^{j}e\cdot(z_{2}-z_{1})$. Here $e(t)=(R(t)-y(t))$ represents the reward prediction error for the last choice and $\eta$ is the learning rate, controlling the speed of adaptation. As the latter learning rule features non-local terms, it can easily be implemented by a network that dynamically adjusts the activities of two modulatory populations of virtual weights $w_{1},w_{2}$ without relying on synaptic updates (see Fig. 1 C). In Fig. 1 we compare the performance of an agent that learns with synaptic plasticity with an agent that performs dynamic adaptation. As we can see in panel 1 F, decreasing the timescales of the learning process, the network with dynamic adaptation achieves maximum reward after seeing only one in-context example, while the performance of the network with synaptic plasticity plateaus at near chance level. This is supported by the psychometric curve shown in Fig. 1 G. Interestingly, the comparison between the two curves in Fig. 1 G closely mirrors findings from a recent study [30]. In that study, some human participants displayed the ability to adjust their behavior after just one example, while others, with impaired cognitive abilities, had to relearn the previous classification rule through trial and error from scratch. In this work, we propose biologically plausible framework for understanding rapid behavioral adaptations in dynamic environments, bridging the gap between biological and artificial intelligence.
Figure 1: Adapting behavior in a context-dependent classification task. A. Overview of the task structure: At each iteration, the agent is presented with a cue card and two guide cards, one matching the cue by color and the other by shape. The agent must select the correct guide card to receive a positive reward. The classification rule switches after 20 trials. B. Schematic representation of the proposed network architecture: synaptic plasticity of physical weights is replaced by dynamic adaptation of virtual weights C. that modulate the network response. D. Example temporal trajectory for the plastic network. Two in-context errors are needed to learn the correct rule. E. Example temporal trajectory for the network with dynamic adaptation: one in-context errors are needed to learn the correct rule. F. Average reward after 1 in-context example as a function of the learning rate $\eta$. Synaptic plasticity with a local delta rule does not go beyond the chance level. Dynamic adaptation with a non-local learning rule can achieve a perfect score. The curve is obtained by averaging over 10 independent sessions of 2000 iterations each. The shaded region represents the standard deviation confidence interval. G. Psychometric curve comparing average reward as a function of in-context examples. At $-1$ we report the average reward before the rule switch. The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
@@
Consider a general learning scenario, where an agent is engaged in an environment $\alpha$ from a family of tasks $\mathcal{A}$. The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f$ from the environment in return. This can be an error, reward, or the target behavior.
@@
In this interpretation, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights. We show that such networks can dynamically implement standard gradient-based algorithms widely used in machine learning. While these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the circuit level. We demonstrate that such circuits capable of dynamic learning can simply arise by tuning the static readout weights of random RNNs with multiplicative interactions, requiring only a limited number of training trajectories from the original learning algorithm applied to a subset of the task family. The construction of our networks follows two distinct phases: a developmental or evolutionary phase, and a subsequent phase where the network leverages this structure for dynamical learning.
@@
Recent research has highlighted a growing interest in understanding modulatory mechanisms and their role in neural computation. It has been observed that gain modulation allows neurons to adjust their response to inputs based on context [24]. This dynamic regulation of neuronal sensitivity plays a critical role in adaptive behavior and cognitive flexibility. Moreover, it has been implicated in various functions, including feedback control [32, 33]) and attention, where it allows for the selective amplification of relevant information. A recent study [34] identified gain modulation as a key neural mechanism underlying hierarchical spatiotemporal prediction and sequence learning, demonstrating that such models can replicate various cortical processing features, including temporal abstraction and space-time receptive fields in the primary visual cortex. Furthermore, gain modulation is closely tied to meta-learning and multitask learning [35, 36], enabling neural networks to efficiently generalize across diverse tasks by leveraging contextual cues.
@@ Results
Figure 4: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). C. Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD). Line: median, shading: 20-th/80-th percentile range. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
Reinforcement learning in a reaching task.
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [39]),
set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility. The position of the agent is encoded separately using 25 input units $\mathbf{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
@@ Discussion
While we do not fully address how the static weights of the networks involved can be learned using biologically plausible learning rules, we demonstrate that these architectures naturally emerge by tuning the readout weights of random networks. These weights could be shaped through evolutionary processes, adapted over slow timescales via reinforcement learning, or optimized using local plasticity rules based on feedback alignment. We argue that this represents a step forward in biological plausibility compared to existing deep-learning models like transformers, gated RNNs, and hyper-networks, which rely on more complex and biologically implausible credit assignment mechanisms such as backpropagation through time to train synaptic weights.

### v5

@@ Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated learning theories to synaptic plasticity. However, such rapid changes are not coherent with the timescales of biological synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical reconfiguration of the network involved, without changing its weights. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks, such as in pyramidal neurons thanks to input segregation and dendritic amplification. We propose a constructive approach to induce in-context learning in an architecture composed of recurrent networks with gain modulation, demonstrating abilities inaccessible to standard networks. In particular, we show that, such architecture can dynamically implement standard gradient-based by encoding weight changes in the activity of another network. We argue that, while these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the level of neural circuits. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. We further validate our approach in a realistic robotic setting, demonstrating dynamic adaptation in a MuJoCo ant navigation task without synaptic updates, showcasing a neuromorphic control paradigm that operates through real-time network reconfiguration. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence, and provides a measurable approach to distinguish between dynamic adaptation and synaptic plasticity, aiding experimental validation.
@@ Introduction
Main
@@
Consider a general learning scenario, where an agent is engaged in an environment $\alpha$ from a family of tasks $\mathcal{A}$. The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f(t)$ from the environment in return. This can be an error, reward, or the target behavior.
@@
In the Supplementary Materials, we show that a non-local weight update rule, implemented through dynamic adaptation, enables rule switching with a single error feedback in a context-dependent classification task, whereas local synaptic plasticity drops to chance level after one error—independently of the timescales of the learning process (Fig. A2). In the dynamic adaptation setting, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights. We show that such networks can dynamically implement standard gradient-based algorithms widely used in machine learning. While these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the circuit level. We demonstrate that such circuits capable of dynamic learning can simply arise by tuning the static readout weights of random RNNs with multiplicative interactions, requiring only a limited number of training trajectories from the original learning algorithm applied to a subset of the task family. The construction of our networks follows two distinct phases: a developmental or evolutionary phase, and a subsequent phase where the network leverages this structure for dynamical learning.
@@
Recent research has highlighted a growing interest in understanding modulatory mechanisms and their role in neural computation. It has been observed that gain modulation allows neurons to adjust their response to inputs based on context [24]. This dynamic regulation of neuronal sensitivity plays a critical role in adaptive behavior and cognitive flexibility. Moreover, it has been implicated in various functions, including feedback control [31, 32] and attention, where it allows for the selective amplification of relevant information. A recent study [33] identified gain modulation as a key neural mechanism underlying hierarchical spatiotemporal prediction and sequence learning, demonstrating that such models can replicate various cortical processing features, including temporal abstraction and space-time receptive fields in the primary visual cortex. Furthermore, gain modulation is closely tied to meta-learning and multitask learning [34, 35], enabling neural networks to efficiently generalize across diverse tasks by leveraging contextual cues.
@@ Results
Figure 3: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. C. Illustration of the task: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. C. Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD). Line: median, shading: 20-th/80-th percentile range. B, D, F. Similar to A, C, E, respectively, but without gain modulation.
Dynamical Reinforcement learning in a reaching task
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [38]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility. The position of the agent is encoded separately using 25 input units $\mathbf{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
@@
Application to a realistic task
We test our framework to a more complex and realistic setting by applying it to a MuJoCo ant environment, demonstrating its scalability to continuous control tasks. The agent, modeled as a quadruped, operates in a high-dimensional control space with 27 environmental state variables and 8 control torques (2 per leg), requiring coordinated control of multiple degrees of freedom.
To manage low-level motor execution, we first trained a network to perform four fundamental motor primitives—moving left, right, forward, and backward. This pretraining step allows higher-level adaptation without direct joint-level optimization (see Appendix for details). Building on this, we use our gain-modulated network as a meta-policy that dynamically selects motor primitives based on the agent’s position. The meta-policy network (orange network Fig. 4A) receives positional inputs and determines the appropriate primitive, while a secondary network (green network Fig. 4A) modulates its behavior through virtual weights, mirroring the architecture used in the grid-based task.
This architecture enables flexible and efficient adaptation: the agent successfully generalizes to target locations not included in the training set (see Fig. 4C), adjusting its movement strategies dynamically without modifying its underlying synaptic structure. This capability is particularly significant for neuromorphic computing and robotics, where adaptive behavior is often constrained by the need for on-chip plasticity. Our results demonstrate that a fixed neuromorphic chip, programmed with a gain-modulated network, can dynamically adjust to new goals without requiring structural modifications. This approach reduces computational overhead and power consumption while enabling efficient, low-power autonomous systems that adapt in real time. By eliminating the need for weight updates at the hardware level, this architecture provides a scalable and robust foundation for adaptive physical agents, paving the way for next-generation neuromorphic robotics.
Figure 4: Dynamical Reinforcement Learning Dark Room A. Overview of the network architecture employed. One GM-network (depicted in green) is responsible for estimating the necessary update of virtual weights through policy gradient, while another GM-reservoir is focused on estimating policy reconfiguration (shown in orange) resulting from changes in virtual weights. B. Illustration of the task, and of the mojoco environment: the agent begins from the center and learns to reach not observable objects positioned at various locations (represented by colored circles). After several trials, the agent achieves precise targeting of the circles. Our networks are pre-trained on a set of target points (red crosses) an then tested on the same (displayed in the left panel) and new positions (shown in the right panel). C. Reward as a function of number of trials (or games), blue: policy gradient, left panel pink: dynamical learning for training positions (ID), right panel pink: dynamical learning for new positions (OOD). Line: median, shading: 20-th/80-th percentile range.
@@ Discussion
Our findings demonstrate that adaptive network architectures can emerge by selecting appropriate readout weights in initially random networks, without requiring explicit backpropagation. These weights can be shaped through biologically plausible mechanisms such as evolutionary processes, slow reinforcement learning, or local plasticity rules relying on feedback alignment. Unlike standard deep-learning models—such as transformers, gated RNNs, and hyper-networks—that depend on non-local credit assignment mechanisms like backpropagation through time, our approach leverages local, dynamically reconfigurable processes, making it more compatible with biological learning principles.
In line with this, our model suggests that dynamic reconfiguration provides a biologically plausible mechanism for rapid adaptation, complementing traditional fast synaptic plasticity. Different neural circuits in the brain may rely on varying balances of dynamic reconfiguration and synaptic plasticity, shaped by their functional roles and evolutionary pressures. In the Supplementary Materials, we provide a proof of concept through a T-maze simulated experiment, where we observe a substantial difference in how contextual information is encoded for synaptic versus adaptive learning. This distinction supports the potential for experimental validation.